OpenStudy (anonymous):

Solve the equation. |2h – 3| = 1 {1, 2} {-1, 2} {-2, 2} @Michele_Laino

OpenStudy (michele_laino):

here we have to distinguish two cases, namely: case 1) 2h-3> or =0 then your equation becomes: \[\Large 2h - 3 = 1,\quad \left( {2h - 3 \geqslant 0} \right)\] please solve for h, what do you get?

OpenStudy (er.mohd.amir):

2h-3=1 and 2h-3=-1 solve these two give ans

OpenStudy (michele_laino):

please, the question is more subtle @Er.Mohd.AMIR

OpenStudy (anonymous):

i pick A am i right

OpenStudy (michele_laino):

I'm sorry, I can not give the direct answer, since I have to respect the Code of Conduct! @jazzy288a

OpenStudy (anonymous):

ik that i was just asking if my answer is right or is it wrong

OpenStudy (michele_laino):

you asked to solve your equation, not to confirm your answer

OpenStudy (michele_laino):

second case is: \[\Large - 2h + 3 = 1,\quad \left( {2h - 3 < 0} \right)\] please solve for h Please, note that not all the algebraic solutions can be the requested solutions